### v2

# Surface Tension of a Topological Phase

Abstract
In interacting topological systems, Landau-like order parameters interplay with the band topology of fermions. The physics of domain formation in such systems can get significantly altered due to the presence of topological fermions. In this work we show that coupling a topological fermionic field to a scalar field can drastically modify the nucleation processes of the scalar field. We find that existence of non-trivial fermionic boundary modes on the nucleating droplets of the scalar field leads to substantial quantum corrections to the surface tension. This leads to an increase in the size of the critical nucleus beyond which unrestricted droplet growth happens. To illustrate the phenomena we devise a minimal model of fermions in a Chern insulating system coupled to a classical Ising field in two spatial dimensions. Using a combination of analytic and numerical methods we conclusively demonstrate that topological phases have a characteristic quantum surface tension. Apart from material systems, our work opens up a host of questions regarding the impact of fermionic topological terms on classical phase transitions and associated criticality.
Introduction: The physics of nucleation, where a droplet having an order parameter value defining the global energy minima grows or shrinks within a metastable phase is one of the defining tenets within the study of phases and phase transitions [1, 2, 3]. These ideas govern a wide range of phenomena spanning condensed matter [4, 5, 6, 7], nuclear physics [8, 9, 10, 11, 12, 13, 14] and high-energy field theories [15, 16, 17, 18, 19, 20, 21]. The conditions on growth and decay of any such nuclei, however, are governed by a competition between surface tension and bulk free-energy density [22, 23, 24]. Microscopic models have been devised to make quantitative predictions and study physical processes such as crystal growth, domain formation and role of thermal excitations and impurity effects [25, 26, 27, 28, 29, 6]. With advances in computational methods and new experimental platforms, many of these ideas have been investigated ranging from material to colloidal systems [30, 31, 32, 33]. While dominantly studied in classical systems, the nucleation physics in presence of quantum fluctuations have been little explored [34]. In particular, the role of topological fermions has not been investigated. Insights from these might be crucial to understand domain formation and nucleation processes in interacting topological insulators where Landau-like order parameters interplay with band topology [35, 36].
Figure 1: Topological droplet: (a) Free-energy density (blue curves) for a two-dimensional classical Ising model coupled to topological fermions in magnetic field $h$ and temperature $T$ where $C$ is Chern number and arrow shows sudden quench. (b) A $C=1$ droplet within a metastable region of $C=-1$. (c) A droplet of size $n$ either grows or shrinks depending on a critical size. This critical size for classical Ising system ($n_{c}^{\rm{cl}}$) increases to $n_{c}^{\rm{Q}}$ when coupled to topological fermions due to droplet edge-modes leading to additional surface tension.
In this work we pose and solve the question – can a topological fermionic field, when coupled to a scalar field, modify the physics of nucleation? We answer this in affirmative, uncovering the role of topological edge-modes in modifying the surface tension and nucleation physics. Such edge-modes are a hallmark of topological phases of matter where they retain their robustness even when the bulk is gapped [37, 38, 39, 40, 41]. While our investigation here is theoretical, such systems of coupled fermionic degrees of freedom with Landau-like scalar order parameters have been of immense interest given their direct applicability to a range of material platforms including those of anomalous quantum Hall systems, spintronics and physics of correlated topological phenomena [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54].
We limit our study to a scalar Ising field in two dimensions which undergoes the standard order-disorder transitions in the magnetic field ($h$) and temperature ($T$) plane (see Fig. 1(a)) where the second order thermal phase transition happens at $T=T_{c}$ [55]. We couple a two dimensional fermionic Chern insulator to the Ising field in such a way that the order parameter of scalar field acts like effective mass for the Dirac fermions. Thus the free-energy minima for $h>0$ and $h<0$ (for $T<T_{c}$) regime also correspond to topological insulating phases (for fermions) albeit with different Chern numbers ($C=\pm 1$). We perform a gedanken experiment where a quench is done from $h<0$ to $h>0$ regime such that the system is in a metastable phase of $C=-1$ with spins down polarized while the global free-energy minima is for $C=+1$ phase with spins up polarized. We now investigate the stability of a bubble of $C=+1$ region created within this metastable phase (see Fig. 1(b)). In absence of any fermionic field it is known that a bubble only beyond a critical size $n_{c}^{\rm{cl}}$ grows, while droplets of size $n<n_{c}^{\rm{cl}}$ shrinks with time. Here, we demonstrate that presence of a topological fermionic field when coupled to a scalar field leads to quantum corrections to surface tension essentially due to topological edge-modes on droplet boundary (see Fig. 1(b)). This quantum surface tension leads to an enhancement of the critical nuclei size to $n_{c}^{\rm{Q}}$ such that $n_{c}^{\rm{Q}}$ is always greater than $n_{c}^{\rm{cl}}$ (see Fig. 1(c)). In fact as we will show such enhancement can be tuned by changing the microscopic parameters of theory, easily achieving $n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}\sim 2$ within experimentally relevant scales.
Model: We consider Bernevig-Hughes-Zhang (BHZ) model [56] of spinless fermions coupled to two-dimensional ferromagnetic classical Ising model on a square lattice, such that the Hamiltonian is
$H=-\kappa\sum_{i}s_{i}{\bf{\Psi}}_{i}^{\dagger}\sigma_{z}~{}{\bf{\Psi}}_{i}-\sum_{\langle ij\rangle}({\bf{\Psi}}_{i}^{\dagger}{\large{\eta}}_{ij}~{}{\bf{\Psi}}_{j}+{\rm{h.c.}})\\ -J\sum_{\langle ij\rangle}s_{i}s_{j}-h\sum_{i}s_{i},$ (1)
where the spin variable for $i$-th unit cell can assume $s_{i}=\pm 1$ and $J>0$. Each unit cell contains $A$ and $B$ sites having staggered masses ±$\pm\kappa s_{i}$. The hopping strengths between nearest neighboring unit cells along $x$-axis and $y$-axis are ${\large{\eta}}_{ij}=\frac{1}{2}(\sigma_{z}+i\sigma_{x})$ and ${\large{\eta}}_{ij}=\frac{1}{2}(\sigma_{z}+i\sigma_{y})$ respectively where $\sigma_{x}$, $\sigma_{y}$, $\sigma_{z}$ are Pauli matrices and ${\bf{\Psi}}_{i}=\begin{pmatrix}c_{iA}&c_{iB}\end{pmatrix}^{\rm{T}}$ with $c_{iA}$ ($c_{iB}$) being annihilation operator of fermion at site $A$ ($B$) of $i$-th unit cell. $\kappa$ is coupling parameter between the fermions and Ising spins. We keep fermionic filling fixed at half. When $\kappa=1$, an up polarized state with magnetization $m=\frac{1}{L^{2}}\sum_{i}\langle s_{i}\rangle=+1$ corresponds to a fermionic Chern insulator with $C=+1$ and a down polarized state ($m=-1$) leads to $C=-1$. Moreover at $T=0$, $\kappa$ itself can be tuned such that when $20<|\kappa|<$ 2 we have a topological phase with $C=\pm 1$ and a trivial phase with $C=0$ when $|\kappa|>2$ [56, 57, 58]. We will consider periodic boundary conditions throughout our work.
[…]
Quantum correction to surface tension: When $\kappa=0$, the fermions and spins are decoupled and the physics of nucleation under the quench protocol is determined by just the classical Ising fields. Thus, change of free-energy in the formation of a circular cluster of size $n$ is
$\Delta F_{\rm{classical}}=-2|h|n+2{\sqrt{\pi n}}\sigma_{\rm{cl}},$ (2)
where $\sigma_{\rm{cl}}>0$ is the surface tension which depends on $J$ and temperature $T$ [2, 27]. When $\kappa=0$, change of free energy $\Delta F_{\rm{classical}}$ is maximum for critical cluster-size
$n_{c}^{\rm{cl}}=\frac{\pi\sigma_{\rm{cl}}^{2}}{4|h|^{2}}.$ (3)
Thus, to reduce the free energy, the cluster grows with time when $n>n_{c}^{\rm{cl}}$ and it shrinks when $n<n_{c}^{\rm{cl}}$, determined by a competition of surface tension with bulk energy-density and a system with a higher surface tension $\sigma_{\rm{cl}}$ would pertain to a larger $n_{c}^{\rm{cl}}$.
Figure 2: Surface tension and critical cluster-size: (a) Change of ground-state energy $\Delta E$ for fermionic system with $\sqrt{n}$ where $n$ is cluster-size. (Inset) Correction to surface tension $\Delta\sigma$ obtained from slope of linear fit (see Eq. (5)) in (a) as a function of $\kappa$, where dotted lines denote quantum critical points $\kappa=-2,0,2$. (b) $\gamma=n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}$ (see Eq. (8)) with $\kappa$ and $J$ at $T=0$, where $n_{c}^{\rm{cl}}$ is critical cluster-size for classical situation ($\kappa=0$). Here, number of unit cells is $L^{2}=4096$. In topological situation ($20<|\kappa|<$ 2), $n_{c}^{\rm{Q}}>n_{c}^{\rm{cl}}$ and relative enhancement $\gamma$ increases with decreasing $J$.
Figure 3: Critical cluster-size and droplet growth: (a) $\nu$ as a function of cluster-size $n$ for classical ($\kappa=0$), topological ($\kappa=1.2$) and trivial ($\kappa=2.5$) situations (see Eq. (9)). $\nu=0$ denotes the critical sizes which are $n_{c}^{\rm{cl}}=145$, $n_{c}^{\rm{Q}}=304$, $n_{c}^{\rm{Q}}=158$ respectively. While in topological situation $n_{c}^{\rm{Q}}>n_{c}^{\rm{cl}}$, in trivial situation $n_{c}^{\rm{Q}}\approx n_{c}^{\rm{cl}}$. (b) Initial and final average local magnetization $\langle s_{x,y}\rangle$ when clusters of sizes, $n=441$ ($n>n_{c}^{\rm{Q}}$), $n=225$ ($n_{c}^{\rm{Q}}>n>n_{c}^{\rm{cl}}$), $n=81$ ($n_{c}^{\rm{cl}}>n$) are evolved. The final time $t=100$ is in units of MC steps/site. Upper (lower) panel shows $\kappa=0$ ($\kappa=1.2$) and $J=0.2$, $T=1/\beta=0.33<T_{c}\approx 0.45$, $h=0.02$, $L^{2}=4096$. Averaging is performed over $10^{3}$ simulations. For classical ($\kappa=0$) situation, the cluster grows when $n>n_{c}^{\rm{cl}}$ but in topological situation ($\kappa=1.2$), growth occurs only when $n>n_{c}^{\rm{Q}}$.
When $\kappa\neq 0$, the Ising field and the fermionic system are coupled thus changing the total free-energy of the system. Under the same quench protocol, total change of free-energy $\Delta F$ is now
$\Delta F=\Delta F_{\rm{classical}}+\Delta E,$ (4)
where $\Delta E$ is the quantum contribution from fermionic field which can again have both surface and bulk contributions. The bulk contribution from the fermionic field is zero since in the nucleating droplet there is a flip of the effective mass term $M_{\text{eff}}\equiv\kappa m$ under which the ground state energy of fermionic system at half-filling does not change. The surface contributions to $\Delta E$ can however be non-zero depending on the character of the fermionic phases as we discuss below.
To estimate $\Delta E$ we compare the ground state energies ($h=0,T=0$) between configurations (i) where a square region of size $L^{2}$ has all $s_{i}=-1$ and another (ii) where an internal region of size $l^{2}$ (where $l=\sqrt{n}$) has $s_{i}=+1$ while the rest of $s_{i}=-1$. Variation of $\Delta E$ shows a characteristic $\sqrt{n}$ dependence for a finite $\kappa$ (see Fig. 2(a)), thus leading to a quantum contribution to surface tension ($\Delta\sigma$[59, 60] which can be estimated from the slope as
$\Delta E=4(\Delta\sigma)\sqrt{n},$ (5)
where $4\sqrt{n}$ is the perimeter of the internal region. Furthermore $\Delta\sigma$ has an interesting dependence on $\kappa$ which we discuss next.
As shown in inset of Fig. 2(a), $\Delta\sigma\rightarrow 0$ when $\kappa\rightarrow 0$ reflecting the decoupled limit. However $\Delta\sigma$ is significantly larger in topological situation ($20<|\kappa|<$ 2) compared to the trivial situation ($|\kappa|>2$). In the topological situation the droplet necessarily hosts robust edge-modes at the boundary due to a change in the Chern number between the cluster and its background. This imposes a reorganization of the energy-spectrum, causing fermions to move up from the bulk energy levels to mid-gap edge modes. Given a bulk gap of $\sim\Delta_{\rm{bg}}$, the edge state contribution to surface tension can be estimated to be
$\Delta\sigma_{\text{edge}}\sim\frac{\Delta_{\rm{bg}}^{2}}{\hbar v_{F}}$ (6)
where $v_{F}$ is the Fermi velocity of the edge modes. This should be contrasted with the trivial situation where no such edge-mode appears resulting in $\Delta\sigma\rightarrow 0$. Thus $\Delta\sigma$ in topological situation is primarily determined by edge-mode energy per unit length ($\Delta{\mathcal{E}}_{\text{edge}}$) of the boundary of droplet. In general, a nucleating domain of one fermionic phase of matter into another, even both being trivially insulating, has a non-zero surface tension contribution due to perturbative reorganization of all bulk states resulting from domain boundary. Such contributions are, however, non-universal and much smaller than those from topological edge modes. A demarcation of such contributions is discussed in supplemental material (SM) [61].
When $\kappa\neq 0$, total change of free energy $\Delta F$ is maximum for the critical circular cluster-size
$n_{c}^{\rm{Q}}=\frac{\pi\left(\sigma_{\rm{cl}}+\Delta\sigma\right)^{2}}{4|h|^{2}}.$ (7)
Thus the enhancement of critical nuclei of classical Ising field, due to its coupling to fermionic field ($\kappa\neq 0$) can be quantified by defining
$\gamma=\frac{n_{c}^{\rm{Q}}}{n_{c}^{\rm{cl}}}=\left(1+\frac{\Delta\sigma}{\sigma_{\rm{cl}}}\right)^{2}.$ (8)
Interestingly, $\gamma$ behaves non-monotonically with $\kappa$, mirroring the behavior of $\Delta\sigma$. The interplay of Ising spin-exchange scale $J$ and edge-mode contribution to $\Delta\sigma$ allows for further tunability of $\gamma$. At $T=0$, surface tension in classical Ising model is $\sigma_{\rm{cl}}=2J$, implying an increase in $\gamma$ with decreasing $J$ in topological situation ($20<|\kappa|<$ 2). For $J\sim 0.2$, $\kappa\sim 1$, one can tune $\gamma\sim 1.5$ which is upto $50\%$ rise in the size of critical nuclei; while in trivial situation ($|\kappa|>2$), $\gamma\sim 1$, implying no significant change in critical cluster-size (see Fig. 2(b)).
While our discussion of topological surface tension has been done in context of model introduced in Eq. (1), the phenomenology holds true in general. For instance, the Hamiltonian here contains, in addition to bare particle-hole symmetry for fermions, a symmetry where the fermionic spectrum remains invariant under spin flip $s_{i}\rightarrow-s_{i}$. This results into the fermionic volume contribution to free energy $\Delta\epsilon=0$ where $\Delta E\propto\Delta\epsilon n$ (see near Eq. (4)). The quench physics then is solely determined by surface term. However, in absence of this symmetry, when $\Delta\epsilon\neq 0$, topological surface tension still dominantly governs the nucleation physics in topological regime albeit with renormalized $n^{\rm{Q}}_{c}$ and $\gamma$ (see SM [61]). To further study if this physics remains stable under thermal fluctuations, we now analyze the system at finite temperatures.
Growth of clusters at finite temperature: At finite $T$, thermal fluctuation reduces the stiffness of domain wall thereby reducing $\sigma_{\rm{cl}}$ from its $T=0$ value ($2J$[27]. It is useful to note we work with $T<T_{c}$ where classical transition temperature $T_{c}\approx 2.269J$ [55] (where Boltzmann constant is set to unity). Moreover, we consider temperatures much smaller than fermionic energy-scales, thus the effect of thermal fluctuation on quantum surface tension can be ignored. Thus with increasing temperature, since effective $\sigma_{\rm{cl}}$ reduces, we expect that $\gamma$ increases [61].
For numerical computation of $n_{c}^{\rm{cl}}$ and $n_{c}^{\rm{Q}}$ at finite $\beta=1/T$, we resort to Monte Carlo (MC) simulation where both the spins and fermions contribute to total energy of any configuration. However, given electronic time-scales are much faster than the spin time-scales, the electronic effects can be effectively included as internal energy corrections to the spin configurations within Metropolis algorithm [62] (see SM [61]). We start from an initial configuration with $n=l^{2}$ up-spins while rest of the spins are down. We then investigate the quantity
$\nu(n)=\frac{P_{+}(n)-P_{-}(n)}{{P_{+}(n)+P_{-}(n)}}$ (9)
where $P_{+}(n)$ and $P_{-}(n)$ count the number of MC steps in which size $n$ of a cluster increases to $(n+1)$ and decreases to $(n-1)$ respectively [6]. Critical cluster-sizes $n_{c}^{\rm{cl}}$ for classical situation ($\kappa=0$) and $n_{c}^{\rm{Q}}$ for both topological ($\kappa=1.2$) and trivial ($\kappa=2.5$) situations are then obtained from when $\nu$ crosses 00 [63], where we find $n_{c}^{\rm{Q}}>n_{c}^{\rm{cl}}$ in topological situation and $n_{c}^{\rm{Q}}\approx n_{c}^{\rm{cl}}$ in trivial situation (see Fig. 3(a)) thus showing the characteristic rise of surface tension only in the topological case.
To observe the growth (shrinkage) of clusters with time $t$, we calculate average local magnetization $\langle s_{x,y}\rangle$ for any unit cell at position ${\textbf{r}}=(x,y)$ as a function of $t$. The behavior of $\langle s_{x,y}\rangle$ for different values of $\kappa$ and $t$ is shown in Fig. 3(b). Our study reveals that when initial size of cluster (having up-spins) $n>n_{c}^{\rm{Q}}$, the cluster grows with time and when $n_{c}^{\rm{cl}}>n$, the cluster shrinks. This behavior remains unchanged in both the classical ($\kappa=0$) and the topological ($\kappa=1.2$) situations. However, when $n_{c}^{\rm{Q}}>n>n_{c}^{\rm{cl}}$, although the cluster grows for $\kappa=0$, it shrinks for $\kappa=1.2$ (see Fig. 3(b)), thus establishing the role of topological surface tension in altering growth (shrinkage) of clusters. Moreover, modifying the quench protocols can further allow tunability in such growth processes (see SM [61]).
Figure 4: Edge-state evolution: Local density of states (LDOS) of edge-modes for initial cluster-sizes (a) $n=441$ ($n>n_{c}^{\rm{Q}}$), (b) $n=225$ ($n<n_{c}^{\rm{Q}}$) at times $t=0,40,80$ (in the units of Monte Carlo steps per site), where $n_{c}^{\rm{Q}}=304$. The parameters chosen here are: $J=0.2$, $T=1/\beta=0.33$ (where $T_{c}\approx 0.45$), $h=0.02$, $\kappa=1.2$. The number of unit cells in square lattice is $L^{2}=4096$. We observe that LDOS of edge-modes captures the signatures of growth (when $n>n_{c}^{\rm{Q}}$) and shrinkage (when $n<n_{c}^{\rm{Q}}$) of clusters for $20<|\kappa|<$ 2.
Edge-state evolution: In order to observe the edge-states in a nucleating droplet, we take a typical spin configuration and solve the fermionic problem at every step in the MC evolution  [61] to obtain the local density of states (LDOS) defined as ${\rm{LDOS}}=\sum_{j,\alpha}|\langle i\alpha|\psi_{j}\rangle|^{2}$ (where $j$ runs over all eigenstates corresponding to single-particle energy-eigenvalues $|\epsilon_{j}|<0.05W$ near the Fermi energy where $W$ is the bandwidth of single-particle spectrum) at $i$-th unit cell with coordinate $(x,y)$ and orbitals $\alpha=A,B$ at various times $t$. Unsurprisingly the edge-state evolution in topological situation ($20<|\kappa|<$ 2) follows the signatures of growth and shrinkage of clusters. We find that edge-modes are indeed localized at the boundary of clusters, however, unlike the configuration averaged circularly symmetric clusters (see Fig. 3(b)) any typical configuration has a rugged domain wall between two different topological regions (see Fig. 4). When $n>n_{c}^{\rm{Q}}$, the boundary of cluster grows with time $t$, while it shrinks and eventually vanishes when $n<n_{c}^{\rm{Q}}$ (see Fig. 4(a,b)). Taken together, all these results show a characteristic surface tension which can be alluded to a topological phase.
Outlook: In this work, we investigate nucleation of a scalar field when coupled to a topological fermionic field. Taking a concrete example of an Ising model coupled to a Chern insulator, we find that fermionic quantum contribution leads to an additional correction to surface tension of a nucleating cluster. Interestingly, the coupling parameter of the two fields can serve as a new tunable parameter for controlling the critical cluster-size in nucleation, thus affecting its growth and shrinkage. Ranging from implications on fundamental aspects of coupled topological field theories, just nucleation physics may also play a role in formation of topological domains such as Chern mosaic structures seen in strongly correlated topological systems [36]. Our work opens up a range of questions regarding physics of nucleation when coupled to higher dimensional topological systems where one can systematically pose the role of such surface physics in modifying critical properties of the classical fields. The imminent breakdown of adiabaticity would result in interplay of quantum and thermal fluctuations leading to intriguing critical theories with characteristic exponents. Furthermore, the role of Goldstone fluctuations when the classical field has continuous symmetry can also be interesting. The study of spinodal decomposition and coarsening dynamics [64, 65] could be important future directions.
Acknowledgements: We acknowledge fruitful discussions with Amit Agarwal, Vijay B. Shenoy, Diptarka Das, Sabyasachi Chakraborty, Soumya Sur, Subrata Pachhal, Rahul Singh. S.M. acknowledges support from PMRF Fellowship, India. AA acknowledges support from IITK Initiation Grant (IITK/PHY/2022010). Numerical calculations were performed on the workstations Wigner and Syahi at IITK.
Supplemental Material to “Surface Tension of a Topological Phase”
I Sources of Surface Tension
In this section we demarcate both the bulk and edge contributions to the surface tension $\Delta\sigma$ for the model discussed in the main text.
I.1 Calculation of edge mode energy (Analytical)
To compute edge-mode energy per unit length for BHZ model, we consider the Hamiltonian
$H=-M\sum_{i}{\bf{\Psi}}_{i}^{\dagger}\sigma_{z}~{}{\bf{\Psi}}_{i}-\sum_{\langle ij\rangle}({\bf{\Psi}}_{i}^{\dagger}{\large{\eta}}_{ij}~{}{\bf{\Psi}}_{j}+{\rm{h.c.}}).$ (S1)
For periodic boundary conditions in both $x$ and $y$ directions, we resort to momentum space where Hamiltonian for $k_{x},k_{y}\in[-\pi,\pi)$ is
$H_{k_{x},k_{y}}=\sin(k_{x})\sigma_{x}+\sin(k_{y})\sigma_{y}+(-M-\cos(k_{x})-\cos(k_{y}))\sigma_{z}.$ (S2)
Thus, the dispersion relation is given by
$E(k_{x},k_{y})=\pm\sqrt{\sin^{2}(k_{x})+\sin^{2}(k_{y})+(M+\cos(k_{x})+\cos(k_{y}))^{2}}.$ (S3)
Now, in a ribbon geometry with periodic (open) boundary condition in $x$ ($y$) direction, $k_{x}$ remains a good quantum number and hosts $x$ dispersing edge states on the top and bottom boundaries in the topological phase ($20<|M|<$ 2). The dispersion in the two situations (open and periodic boundary conditions) and the local density of states (LDOS) of the edge-states are shown in Fig. S1(a-c).
Figure S1: Spectrum and edge modes in BHZ model: (a) Single-particle energy spectrum as a function of $k_{x}$ for BHZ model with $M=1.0$ for periodic boundary in $x$ direction and open boundary in $y$ direction, (b) single-particle energy spectrum as a function of $k_{x}$ for BHZ model with $M=1.0$ for periodic boundary in both $x$ and $y$ directions, (c) Local density of states (LDOS) of edge-modes at unit cell with coordinate $(x,y)$ for BHZ model with periodic boundary in $x$ direction and open boundary in $y$ direction for $L=32$, $M=1.0$.
The difference in the ground-state energy for a half-filled system between the two situations (a) and (b) determines the mean edge-mode energy when the system is deep in the topological phase ($0\ll|M|\ll 2$). For instance, at $M=1$ an edge state wavefunction that decays exponentially in the $y$ direction has an effective energy $E_{\rm{edge}}(k_{x})\sim-\sin(k_{x})$ near $k_{x}=\pi$ (see Fig. S1(a)). However, such an edge state can only be defined appropriately until it mixes with the bulk modes. The valence bulk band maxima $E_{\rm{valence}}(k_{x})$ is determined by $k_{y}=\pi$ mode when $20<M<$ 2 (similarly by $k_{y}=0$ mode when $-2<M<0$). Thus the range of $k_{x}$ until which the edge states survive is given by $k^{c}_{x}=\cos^{-1}(-M+1)=\pm\frac{\pi}{2}$ for $M=1$. The edge state energy is thus given by the difference
[…]
For a $L\times L$ square lattice with periodic boundary condition in $x$ direction and open boundary condition in $y$ direction, the edge length $=2L$ and thus the edge-mode energy per unit length is
[…]
which can be evaluated exactly and is $=\frac{(\pi-2)}{4\pi}$ for $M=1$. For a general linearly dispersing edge state where the dispersion is given by
$E_{\text{edge}}(k_{x})\sim\hbar v_{F}k_{x}.$ (S6)
For the bulk band-gap (energy difference between the valence band maximum to conduction band minimum) $\Delta_{\rm{bg}}$, $k_{x}^{c}$ can be estimated as $\sim\frac{\Delta_{\rm{bg}}}{2\hbar v_{F}}$, thus
[…]
In the present Hamiltonian, since $v_{F}\sim 1$, and thus the surface tension arising from the edge is expected to vary as the square of the bulk gap. In the next section, we compare this analytic estimate with the numerical results of edge and bulk contribution to the surface tension.
Figure S2: Contribution of edge-modes and bulk states to surface tension: (a) Surface tension $\Delta\sigma$, contributions of bulk states $\Delta\sigma_{\rm{bulk}}$ and edge-modes $\Delta\sigma_{\rm{edge}}$ to surface tension as a function of $\kappa$ where $\Delta\sigma_{\rm{bulk}}$ and $\Delta\sigma_{\rm{edge}}$ are obtained using Eq. (S11a) and Eq. (S11b) respectively. (b) Difference of energy $\Delta E_{j}$ of $j$-th single-particle level (see Eq. (S8)) below the Fermi energy in topological situation ($\kappa=1.2$) when the size of the droplet is $n=l^{2}$. (c) $\Delta E_{j}$ as a function of $j$ in trivial situation ($\kappa=2.5$) when $l=16$. (d) $\Delta\sigma_{\rm{edge}}$ and analytical estimate of $\Delta{\mathcal{E}}_{\rm{edge}}$ (see Eq. (S7)) as a function of $\kappa\equiv M$. In all plots, number of unit cells in square lattice is $L^{2}=4096$.
I.2 Contribution of bulk and edge states (Numerical)
In order to isolate the contributions to the surface tension due to the bulk and edge states separately, we devise the following numerical scheme. We compare the single particle energy eigenvalues between the two cases (i) a uniform square region of size $L^{2}$ with all $s_{i}=-1$ (i.e. before quench), (ii) the case where within the region of size $L^{2}$, a square-shaped droplet of size $n=l^{2}$ with $s_{i}=1$ is now formed (i.e. after quench). Each of these systems has $2L^{2}$ single particle energy eigenvalues which can be enumerated and indexed in ascending order from $j=1,\ldots,2L^{2}$. The difference in the single particle eigenvalues between the two cases is
$\Delta E_{j}=E_{j}(\text{after quench})-E_{j}(\text{before quench})$ (S8)
and their additive sum until the Fermi energy ($j=L^{2}$) is the essential contribution to surface tension $\Delta\sigma$ at temperature $T=0$ which was evaluated in the main text using
$\Delta E=\sum_{j=1}^{L^{2}}\Delta E_{j}=4\Delta\sigma\sqrt{n}.$ (S9)
Given we know that in the topological phase, $l$ eigenstates below the Fermi energy form the edge state manifold, the contribution to $\Delta\sigma$ from the bulk and edge can be separately estimated as
$\Delta E=\Delta E_{\text{bulk}}+\Delta E_{\text{edge}},$ (S10)
where
$\Delta E_{\rm{bulk}}=\sum_{j=1}^{L^{2}-l}\Delta E_{j}=4\Delta\sigma_{\rm{bulk}}\sqrt{n},$ (S11a)
$\Delta E_{\text{edge}}=\sum_{j=L^{2}-l+1}^{L^{2}}\Delta E_{j}=4\Delta\sigma_{\text{edge}}\sqrt{n}.$ (S11b)
Thus as a function of $\kappa$, i.e. the coupling between the spin and the fermion, the complete surface tension $\Delta\sigma$ can be separately decomposed into its bulk and edge contributions where
$\Delta\sigma=\Delta\sigma_{\text{bulk}}+\Delta\sigma_{\text{edge}}.$ (S12)
While in the main text, the variation of $\Delta\sigma$ was shown as a function of $\kappa$, in Fig. S2(a) we show the individual contributions from both the bulk and edge states. Deep in the topological regime $\kappa\sim 1$, one finds that the contribution to $\Delta\sigma$ is dominantly from the edge states. This is reflected even in the $\Delta E_{j}$, where we show its variation as a function of the eigenvalue index $j$ for $\kappa=1.2$, which shows that only $\sim l$ states near the Fermi energy have significant $\Delta E_{j}$ (see Fig. S2(b)). In contrast, when $\kappa=2.5$ such that the system is in a trivial phase (no edge states on the droplet boundary), the sole contribution to $\Delta\sigma$ is due to the bulk states. The behavior of $\Delta E_{j}$ is non-universal and fluctuates over all the bulk states in the trivial phase (see Fig. S2(c)).
A comparison of the $\Delta\sigma_{\text{edge}}$ with the analytic estimate $\Delta{\mathcal{E}}_{\rm{edge}}$ from the last section (where we set $\hbar=1$) is shown in Fig. S2(d), illustrating the topological origin of the surface tension in the regime $20<|\kappa|<$ 2 which is universal and applicable to any topological system with a boundary manifold. Any contribution to $\Delta\sigma$ in the trivial phase $|\kappa|>2$ is due to the bulk states, which is model dependent and non-universal as discussed in the section below.
II Non-Universal Surface Tension of Trivial Insulators
In this section, we discuss how the contribution to $\Delta\sigma_{\text{bulk}}$ is non-universal and model dependent using two different Hamiltonian systems.
II.1 Trivial phases in BHZ model
Figure S3: Surface tension for various filling fractions in trivial situation: (a) Difference of energy $\Delta E_{j}$ of $j$-th single-particle level (see Eq. (S8)) after and before the quench in trivial situation ($\kappa=4.5$) when there is a square shaped droplet of size $n=l^{2}$ (where $l=16$) with $s_{i}=1$, while the rest of $s_{i}=-1$. (b) Surface tension $\Delta\sigma_{\rm{bulk}}^{f}$ as a function of filling fraction $f$ for various values of $\kappa$ in trivial situation of BHZ model. (c) $\Delta E_{j}$ as a function of $j$ in trivial insulator with $\kappa^{\prime}=4.5$ (see Hamiltonian in Eq. (S14)) when there is a square shaped droplet of size $n=l^{2}$ (where $l=16$). (d) $\Delta\sigma_{\rm{bulk}}^{f}$ as a function of filling fraction $f$ for various values of $\kappa^{\prime}$ in trivial insulator . In all plots, number of unit cells is $L^{2}=4096$.
We first analyze the variation of the surface tension in trivial situation ($|\kappa|>2$) of the model discussed in the main text, where the $\Delta E_{j}$ for $\kappa=2.5$ was shown in Fig. S2(c). In Fig. S3(a) we show $\Delta E_{j}$ for $\kappa=4.5$, again reflecting the non-universal fluctuation of $\Delta E_{j}$ about zero. Furthermore, we find that both the value and the sign of $\Delta\sigma_{\text{bulk}}$ can be filling ($f$) dependent. In Fig. S3(b), we plot $\Delta\sigma_{\text{bulk}}^{f}$ using
$\Delta E(f)=\sum_{j=1}^{j_{\rm{max}}}\Delta E_{j}=4(\Delta\sigma^{f}_{\text{bulk}})\sqrt{n},$ (S13)
where $j_{\rm{max}}=2L^{2}f$ for different values of $\kappa=2.5,~{}3.5,~{}4.5$ all in the trivial phase of the model discussed in the main text. We find that $\Delta\sigma^{f}_{\text{bulk}}$ switches sign as $f$ is varied. While in this model all the bulk band gap closings are Dirac like - we investigate $\Delta\sigma^{f}_{\text{bulk}}$ in another model below where the bulk gap opening is quadratic.
II.2 Quadratic trivial band insulator
To demonstrate the non-universality of $\Delta\sigma_{\text{bulk}}$ in absence of any edge-mode, we consider another model represented by the Hamiltonian
$H=-\kappa^{\prime}\sum_{i}s_{i}{\bf{\Psi}}_{i}^{\dagger}\sigma_{z}~{}{\bf{\Psi}}_{i}-\frac{1}{2}\sum_{\langle ij\rangle}({\bf{\Psi}}_{i}^{\dagger}\sigma_{z}~{}{\bf{\Psi}}_{j}+{\rm{h.c.}})-J\sum_{\langle ij\rangle}s_{i}s_{j}-h\sum_{i}s_{i},$ (S14)
where ±$\pm\kappa^{\prime}s_{i}$ is the staggered mass of $A$ and $B$ sites of $i$-th unit cell and the hopping is restricted only between $A$-$A$ and $B$-$B$ sites of the nearest neighboring unit cells. For uniform $s_{i}=1$, we assume $M^{\prime}\equiv\kappa^{\prime}$ and the Hamiltonian can be recast in momentum space for $k_{x},k_{y}\in[-\pi,\pi)$ as
$H_{k_{x},k_{y}}=(-M^{\prime}-\cos(k_{x})-\cos(k_{y}))\sigma_{z}.$ (S15)
Thus, the bulk energy-gap is zero for $|M^{\prime}|\leq 2$ where the system behaves as a metal, while for $|M^{\prime}|>2$ the bulk energy-gap is finite and the system is a trivial insulator. Just like the model presented in the main text the spectrum has a $M^{\prime}\to-M^{\prime}$ symmetry but at $M^{\prime}=\pm 2$ the metallic phase gives way to an insulator at half-filling via quadratically dispersing band. Importantly, there is no topological phase in this model. We now consider the same quench protocol as performed for BHZ model discussed in the main text. For a trivial insulator ($|\kappa^{\prime}|\gg 2$), we find that $\Delta E_{j}$ for all the bulk states shows non-universal fluctuation about zero (see Fig. S3(c)). Similar to the trivial phase of BHZ model, here also $\Delta\sigma_{\rm{bulk}}^{f}$ shows non-universal dependence on filling $f$ and $\Delta\sigma_{\rm{bulk}}^{f}<0$ at half-filling (see Fig. S3(d)).
Therefore, the surface tension $\Delta\sigma_{\rm{bulk}}$ in trivial phase at half-filling (i.e. ground state) can be either positive or negative which is decided by the random contributions from individual single-particle energy-levels, rendering the behavior of surface tension in trivial situation non-universal.
III Role of Symmetries and Topological Surface Tension
In this section, we demonstrate the role of symmetries in affecting the surface tension in topological phases. The model discussed in the main text has a set of intricate symmetries such that the contribution of the surface tension can be distilled. These are enumerated below:
  1. 1.
    Spin-fermion flip symmetry: The energy eigenvalues $E_{j}$ of bulk spectrum are independent of sign of $\kappa$, i.e. $E_{j}(\kappa)=E_{j}(-\kappa)$. This also reflects into a spin-flip symmetry as $E_{j}(\kappa,\{s_{i}\})=E_{j}(\kappa,\{-s_{i}\})$. This guarantees that under the quench protocol, the bulk energy density contribution $\Delta\epsilon$ to free energy equals zero. Microscopically this corresponds to a symmetry operation $s_{i}\rightarrow-s_{i},{\bf{\Psi}}_{i}\rightarrow(-1)^{i_{x}+i_{y}}\sigma_{x}{\bf{\Psi}}_{i}\otimes K$ where $K$ is the complex conjugation and $(i_{x},i_{y})$ is the position of $i$-th unit cell.
  2. 2.
    Particle-hole (fermion) symmetry: Under this operation ${\bf{\Psi}}_{i}\rightarrow\sigma_{x}{\bf{\Psi^{\dagger}}}_{i}^{T}$, $s_{i}\rightarrow s_{i}$, $H\rightarrow H$, rendering the fermionic Hamiltonian into the symmetry class D.
Note that the fermion Hamiltonian does not have microscopic time-reversal symmetry - which results in stabilization of finite Chern number phases. In the following, we examine the role of breaking each of these symmetries.
III.1 Role of breaking spin-fermion flip symmetry (class D model)
Figure S4: Effect of hopping between next nearest neighbors: (a) Chern number $C$ as a function of mass parameter $M$, (b) change in energy density $\Delta\epsilon$ as a function of $\kappa$, (c) additional surface tension $\Delta\sigma$ as a function of $\kappa$, (d) enhancement ratio $\gamma=n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}$ of critical cluster size as a function of $\kappa$ when the hopping between next nearest neighboring unit cells is absent ($t_{2}=0.0$) and present ($t_{2}=1.0$). In all plots, $J=0.2$, $h=1.0$, $T=0$, $L^{2}=4096$.
In the previous section we had pointed out that the Hamiltonian in the main text has a symmetry where the sign of the spin does not change the single particle energy eigenvalues. In order to break this symmetry (while retaining particle-hole symmetry), we introduce the next-nearest neighbor hopping $t_{2}$ such that the complete spin-fermion hopping Hamiltonian becomes
$H=-\kappa\sum_{i}s_{i}{\bf{\Psi}}_{i}^{\dagger}\sigma_{z}~{}{\bf{\Psi}}_{i}-\sum_{\langle ij\rangle}({\bf{\Psi}}_{i}^{\dagger}{\large{\eta}}_{ij}~{}{\bf{\Psi}}_{j}+{\rm{h.c.}})-t_{2}\sum_{\langle\langle ij\rangle\rangle}({\bf{\Psi}}_{i}^{\dagger}{\large{\eta}}^{\prime\prime}_{ij}~{}{\bf{\Psi}}_{j}+{\rm{h.c.}})-J\sum_{\langle ij\rangle}s_{i}s_{j}-h\sum_{i}s_{i},$ (S16)
where $t_{2}$ is the strength of the hopping between the next nearest neighboring unit cells and $\eta^{\prime\prime}_{ij}=\frac{1}{2}(\sigma_{z}+i\cos(\theta_{ij})\sigma_{x}+i\sin(\theta_{ij})\sigma_{y})$, the angle $\theta_{ij}$ between the $x$-axis and the bond connecting the next nearest neighboring unit cells can assume the values $\theta_{ij}=\pi/4$ and $\theta_{ij}=3\pi/4$. For uniform $s_{i}$ throughout the system, we assume $M\equiv\kappa s_{i}$ and the Hamiltonian in momentum space for $k_{x},k_{y}\in[-\pi,\pi)$ is written as
$H_{k_{x},k_{y}}$ $=d_{1}(k_{x},k_{y})\sigma_{x}+d_{2}(k_{x},k_{y})\sigma_{y}+d_{3}(k_{x},k_{y})\sigma_{z},$
$d_{1}(k_{x},k_{y})$ $=\sin(k_{x})-\sqrt{2}t_{2}\cos(k_{x})\sin(k_{y}),$
$d_{2}(k_{x},k_{y})$ $=\sin(k_{y})-\sqrt{2}t_{2}\cos(k_{y})\sin(k_{x}),$
$d_{3}(k_{x},k_{y})$ $=-M-\cos(k_{x})-\cos(k_{y})-2t_{2}\cos(k_{x})\cos(k_{y}).$ (S17)
The Hamiltonian satisfies particle-hole symmetry even when $t_{2}\neq 0$, which follows from
$\sigma_{x}H^{*}_{k_{x},k_{y}}\sigma_{x}=-H_{-k_{x},-k_{y}}.$ (S18)
For uniform $s_{i}$, the system can host topological phases having Chern numbers $C=\pm 2,\pm 1$ depending on the value of mass parameter $M$ when $t_{2}\neq 0$ (see Fig. S4(a)). Interestingly, the ground state energy of the fermionic system, which is given by
[…]
where $\text{where~{}}E_{k_{x},k_{y}}$ $=-\sqrt{d_{1}^{2}+d_{2}^{2}+d_{3}^{2}}~{},$ (S19)
is not symmetric under $M\to-M$, i.e., $E_{g}(M)\neq E_{g}(-M)$ when $t_{2}\neq 0$. Thus, the sudden quench from $h<0$ to $h>0$ causes an additional correction to bulk energy density $\Delta\epsilon$ along with the correction to surface tension $\Delta\sigma$ when $t_{2}\neq 0$, such that the change in the energy in the fermionic system for a formation of a square-shaped droplet of size $n$ is
$\Delta E=(\Delta\epsilon)n+4\sqrt{n}(\Delta\sigma).$ (S20)
We show the variation of $\Delta\epsilon$ and $\Delta\sigma$ in Fig. S4(b,c), where we observe that unlike the situation with $t_{2}=0$, $\Delta\epsilon\neq 0$ for $\kappa\neq 0$ when $t_{2}\neq 0$. Furthermore, in the topological situation $20<|\kappa|<$ 2, surface tension $\Delta\sigma$ for $t_{2}\neq 0$ is greater than that for $t_{2}=0$. This can be understood as follows: when $t_{2}\neq 0$ and $20<\kappa<$ 2, a droplet of $s_{i}=1$ corresponds to Chern number $C=2$, while its background with $s_{i}=-1$ has $C=-1$. Thus, the difference in Chern numbers ($|\Delta C|=3$) between the droplet and its background for $t_{2}\neq 0$ is greater than $|\Delta C|=2$ for $t_{2}=0$. An increased $|\Delta C|$ in topological situation causes greater $\Delta\sigma$ since the number of edge modes increases, increasing the effective $\Delta{\mathcal{E}}_{\text{edge}}$. This results in a higher surface tension for $t_{2}\neq 0$ compared to that for $t_{2}=0$. For $-2<\kappa<0$, $\Delta\sigma$ shows similar behavior where a droplet of $s_{i}=1$ has $C=-1$ and its background with $s_{i}=-1$ has $C=2$. In presence of non-zero $\Delta\epsilon$ and $\Delta\sigma$, the critical cluster size for $t_{2}\neq 0$ modifies to
$n_{c}^{\rm{Q}}=\frac{\pi(\sigma_{\rm{cl}}+\Delta\sigma)^{2}}{(2|h|-\Delta\epsilon)^{2}},$ (S21)
resulting in an enhancement ratio
$\gamma=\frac{n_{c}^{\rm{Q}}}{n_{c}^{\rm{cl}}}=\frac{(1+(\Delta\sigma/\sigma_{\rm{cl}}))^{2}}{(1-(\Delta\epsilon/(2|h|)))^{2}}.$ (S22)
We find that in topological situation, $\gamma$ can be greater than $1$ for a suitable choice of $J$ and $h$, indicating an enhancement of critical cluster size (see Fig. S4(d)).
III.2 Role of breaking particle-hole symmetry (class A model)
Figure S5: Effect of particle-hole symmetry breaking hopping term: (a) Chern number $C$ as a function of mass parameter $M$, (b) change in energy density $\Delta\epsilon$ as a function of $\kappa$, (c) additional surface tension $\Delta\sigma$ as a function of $\kappa$, (d) enhancement ratio $\gamma=n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}$ of critical cluster size as a function of $\kappa$ when particle-hole symmetry breaking hopping term is absent ($\lambda=0.0$) and present ($\lambda=0.6$). Dotted lines denote the critical points $\kappa=0,\pm 2$ when $\lambda=0$ and $\kappa=0,\pm 1.6$ when $\lambda=0.6$. In all plots, $J=0.2$, $T=0$, $L^{2}=4096$.
To explore the effect of the hopping term that breaks particle-hole symmetry in BHZ model, we now consider the Hamiltonian
$H=-\kappa\sum_{i}s_{i}{\bf{\Psi}}_{i}^{\dagger}\sigma_{z}~{}{\bf{\Psi}}_{i}-\sum_{\langle ij\rangle}({\bf{\Psi}}_{i}^{\dagger}{\large{\eta}}_{ij}~{}{\bf{\Psi}}_{j}+{\rm{h.c.}})-\lambda\sum_{i}{\bf{\Psi}}_{i}^{\dagger}(\sigma_{x}+\sigma_{y})~{}{\bf{\Psi}}_{i}-J\sum_{\langle ij\rangle}s_{i}s_{j}-h\sum_{i}s_{i},$ (S23)
where $\lambda$ is the hopping parameter between $A$ and $B$ sites within the same unit cell which breaks the particle-hole symmetry. For uniform $s_{i}$ throughout the system, we assume $M\equiv\kappa s_{i}$ and the Hamiltonian in momentum space for $k_{x},k_{y}\in[-\pi,\pi)$ turns out to be
$H_{k_{x},k_{y}}=d^{\prime}_{1}(k_{x},k_{y})\sigma_{x}$ $+d^{\prime}_{2}(k_{x},k_{y})\sigma_{y}+d^{\prime}_{3}(k_{x},k_{y})\sigma_{z},$
$d^{\prime}_{1}(k_{x},k_{y})$ $=(\sin(k_{x})-\lambda),$
$d^{\prime}_{2}(k_{x},k_{y})$ $=(\sin(k_{y})-\lambda),$
$d^{\prime}_{3}(k_{x},k_{y})$ $=(-M-\cos(k_{x})-\cos(k_{y})).$ (S24)
For $\lambda\neq 0$, the particle-hole symmetry of the Hamiltonian is broken, i.e., $\sigma_{x}H^{*}_{k_{x},k_{y}}\sigma_{x}\neq-H_{-k_{x},-k_{y}}$. When $\lambda\neq 0$, the system with uniform $s_{i}$ hosts topological phases with $C=\pm 1$ for $0<|M|<2\sqrt{1-\lambda^{2}}$ and trivial phase ($C=0$) for $|M|>2\sqrt{1-\lambda^{2}}$ (see Fig. S5(a)). Even in absence of particle-hole symmetry ($\lambda\neq 0$), the ground state energy of the fermionic system with uniform $s_{i}$ remains symmetric under $M\to-M$, i.e., $E_{g}(M)=E_{g}(-M)$. Thus, the sudden quench from $h<0$ to $h>0$ causes no additional correction to bulk energy density ($\Delta\epsilon\to 0$), while there is a substantial correction to surface tension $\Delta\sigma>0$ in topological situation (see Fig. S5(b,c)). Therefore, the enhancement ratio $\gamma=n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}=(1+(\Delta\sigma/\sigma_{\rm{cl}}))^{2}$ for $\lambda\neq 0$ shows similar behavior as for $\lambda=0$ (see Fig. S5(d)) and establishes the increase in critical cluster-size even when particle-hole symmetry is absent.
IV Role of Thermal Fluctuations
In this section we discuss the effects of thermal fluctuations on the surface tension and critical droplet size both for the fermions and for the spin degrees of freedom.
IV.1 Fermionic fluctuations
Figure S6: Surface tension and critical cluster-size at finite temperature: (a) Quantum correction to surface tension $\Delta\sigma$ as a function of $\kappa$ at temperatures $T=0$ and $T=1/\beta=0.33$, (b) Surface tension $\sigma_{\rm{cl}}$ of classical Ising model ($\kappa=0$) as a function of $T$, (c) $\gamma=n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}=(1+(\Delta\sigma/\sigma_{\rm{cl}}))^{2}$ as a function of $\kappa$ at temperatures $T=0$ and $T=1/\beta=0.33$, where $n_{c}^{\rm{cl}}$ is the critical cluster-size for classical Ising model ($\kappa=0$). For all the plots, number of unit cells is $L^{2}=4096$. The parameters chosen are: $J=0.2$ (where $T_{c}\approx 0.45$) and $h=0.02$. Here, classical contribution to surface tension $\sigma_{\rm{cl}}=2J$ at $T=0$. At finite $T$ in (b), $n_{c}^{\rm{cl}}$ is evaluated using Monte Carlo simulation for classical Ising model ($\kappa=0$) via Eq. (S26) and $\sigma_{\rm{cl}}$ is then obtained from $n_{c}^{\rm{cl}}$ using Eq. (S27). In (c), $n_{c}^{\rm{cl}}=145$ for $T=1/\beta=0.33$.
In order to distill the role of thermal fluctuation on fermions we freeze the spin configurations before and after quench to where all $s_{i}=-1$ and where a region of size $n=l^{2}$ has $s_{i}=1$. However fermions are now assumed to equilibrate to a finite temperature $T$. The correction to surface tension at a finite temperature $\Delta\sigma(T)$ is again computed from $\Delta E(T)=4\Delta\sigma(T)\sqrt{n}$, where the change in energy $\Delta E(T)$ for the fermionic problem at temperature $T=1/\beta$ (where Boltzmann constant is set to unity) is given by
$\Delta E(T)=\frac{\rm{Tr}(e^{-\beta H_{\rm{final}}}~{}H_{\rm{final}})}{\rm{Tr}(e^{-\beta H_{\rm{final}}})}-\frac{\rm{Tr}(e^{-\beta H_{\rm{initial}}}~{}H_{\rm{initial}})}{\rm{Tr}(e^{-\beta H_{\rm{initial}}})}.$ (S25)
Here, $H_{\rm{initial}}$ is the Hamiltonian for fermionic problem before the quench where all $s_{i}=-1$ and $H_{\rm{final}}$ is the Hamiltonian after the quench where there is a square-shaped cluster of size $n=l^{2}$ having $s_{i}=1$ within the region having $s_{i}=-1$. We find that at a finite temperature $T$, the correction to the surface tension $\Delta\sigma(T)$ does not change significantly from its zero temperature ($T=0$) value in both the topological ($20<|\kappa|<$ 2) and trivial ($|\kappa|>2$) situations, as evident from Fig. S6(a). This is expected since the thermal fluctuations at $T<T_{c}$ are perturbative compared to the fermionic bandwidth.
IV.2 Entropic contributions to spin fluctuations
In order to estimate the role of thermal fluctuations on $\sigma_{\rm{cl}}$ which equals $2J$ at $T=0$, we perform the Monte Carlo simulations to evaluate the critical cluster-size $n_{c}^{\rm{cl}}$ at $\kappa=0$ via
$\nu(n_{c}^{\rm{cl}})=\frac{P_{+}(n_{c}^{\rm{cl}})-P_{-}(n_{c}^{\rm{cl}})}{P_{+}(n_{c}^{\rm{cl}})+P_{-}(n_{c}^{\rm{cl}})}=0$ (S26)
and in turn estimate $\sigma_{\rm{cl}}(T)$ using
$\sigma_{\rm{cl}}(T)=2|h|\sqrt{\frac{n_{c}^{\rm{cl}}}{\pi}},$ (S27)
since any correction to $\sigma_{\rm{cl}}(T)$ for $\kappa=0$ is only due to thermal spin fluctuations and has no fermionic contribution. As shown in Fig. S6(b), finite temperature $T$ reduces $\sigma_{\rm{cl}}$ due to entropic contributions coming from the spin fluctuations in the nucleating region. Given the reduction in the effective surface tension, this illustrates that the spin-fluctuations on the boundary of the nucleating region effectively reduces the stiffness of the domain wall. This is consistent with the analytic estimate of the same (see Ref. [27]).
Interestingly, the effect of thermal fluctuations on fermionic and spin together lead to an increase in the relative enhancement $\gamma=n_{c}^{\rm{Q}}/n_{c}^{\rm{cl}}=(1+(\Delta\sigma/\sigma_{\rm{cl}}))^{2}$ due to an effective reduction of the $\sigma_{\rm{cl}}$. This enhancement of $\gamma$ at finite $T$ is more dominant in the topological situation ($20<|\kappa|<$ 2) compared to the trivial situation ($|\kappa|>2$), given the fermionic contribution $\Delta\sigma$ to surface tension is significant in the topological regime (see Fig. S6(c)).
V Monte Carlo Simulation Details
In this section, we discuss the details of Monte Carlo simulation employed to explore the time evolution of the coupled spin-fermion system.
Figure S7: Effective Monte Carlo simulation: (a,b) Local magnetization $s_{x,y}$ for unit cell at position $(x,y)$ where initial cluster-size with up-spins is $n=441$ ($n>n_{c}^{\rm{Q}}$) at times $t=0,20,40$ (in units of Monte Carlo steps per site) using (a) coupled spin-fermion Monte Carlo simulation (Eq. (S29) and Eq. (S30)) and (b) effective Monte Carlo simulation (Eq. (S31) and Eq. (S32)). The parameters chosen are: $\kappa=1.2$, $J=0.2$, $T=1/\beta=0.33$ (where $T_{c}\approx 0.45$), $h=0.02$. (c,d) Change in ground state energy $\Delta E_{i}^{\rm{Q}}$ (see Eq. (S29)) and its comparison with $\Delta E_{i}^{\rm{Q,eff}}$ (see Eq. (S32)) for fermionic system when $i$-th spin is flipped from (c) $s_{i}=1$ to $s_{i}=-1$, (d) $s_{i}=-1$ to $s_{i}=1$. Here, $N_{i}=\sum_{j\in{\mathcal{N}}_{i}}s_{j}$ where $j$ runs over all the four nearest neighbors (${\mathcal{N}}_{i}$) of $i$-th spin. We find that $\Delta E_{i}^{\rm{Q}}\approx\Delta E_{i}^{\rm{Q,eff}}$. Here, we have chosen $\kappa=1.2$ and number of unit cells in square lattice $L^{2}=4096$.
V.1 Coupled spin-fermion Monte Carlo dynamics
In Monte Carlo (MC) simulation for the classical Ising model ($\kappa=0$), a random spin $s_{i}$ is chosen and the energy $\Delta E_{i}$ required to flip the spin (i.e., $s_{i}\to-s_{i}$) is calculated as:
$\Delta E_{i}=2Js_{i}\sum_{j\in{\mathcal{N}}_{i}}s_{j}+2hs_{i},$ (S28)
where $j$ runs over all the nearest neighbors (${\mathcal{N}}_{i}$) of $i$-th spin. Now, if $\Delta E_{i}<0$ or if $p_{i}\leq\exp(-\beta\Delta E_{i})$ for a randomly chosen $p_{i}\in[0,1]$ where $\beta=1/T$, the spin flip is accepted and the $i$-th spin becomes $-s_{i}$. If none of these conditions is satisfied, the spin-flip is rejected and $i$-th spin remains $s_{i}$.
When classical Ising field is coupled to fermionic field ($\kappa\neq 0$), we solve the fermionic problem to compute $\Delta E_{i}$ at each step of Monte Carlo simulation. In this method, we propose a spin flip $s_{i}\to-s_{i}$ at a randomly chosen $i$-th unit cell at each Monte Carlo step. We then explicitly calculate the difference $\Delta E_{i}^{\rm{Q}}$ in the ground state energies between the configurations after and before the proposed $i$-th spin-flip for the fermionic problem as:
$\Delta E_{i}^{\rm{Q}}=E^{\rm{Q}}\left(s_{1},...,-s_{i},...,s_{L^{2}}\right)-E^{\rm{Q}}\left(s_{1},...,s_{i},...,s_{L^{2}}\right),$ (S29)
where $E^{\rm{Q}}(s_{1},...,s_{i},...,s_{L^{2}})$ denotes the ground state energy for the fermionic system with spin-configuration $s_{1},...,s_{i},...,s_{L^{2}}$. Thus, the total energy required for the $i$-th spin-flip is given by:
$\Delta E_{i}=2Js_{i}\sum_{j\in{\mathcal{N}}_{i}}s_{j}+2hs_{i}+\Delta E_{i}^{\rm{Q}}.$ (S30)
Eq. (S30) can be used to check the acceptance or rejection of the spin-flip at each Monte Carlo step when $\kappa\neq 0$. Under this protocol, the spin fluctuations and their effect on fermions are exactly accounted for. An exact MC simulation for $\kappa=1.2$ and initial droplet size $n=441$ is shown in Fig. S7(a), illustrating the growth of the droplet (as $n>n_{c}^{\rm{Q}}$). This method, however, is numerically expensive, since at every MC step, a matrix diagonalization for the fermionic system needs to be performed. In the next section, we show how this step can be simplified by devising an effective Monte Carlo method.
V.2 Effective Monte Carlo simulation and validity
In order to simplify the exact MC method as discussed in the previous section, we here show that we can also calculate $\Delta E_{i}$ approximately for a spin-flip ($s_{i}\to-s_{i}$) when $\kappa\neq 0$ from the quantum correction to surface tension $\Delta\sigma$ as:
$\Delta E_{i}=2Js_{i}\sum_{j\in{\mathcal{N}}_{i}}s_{j}+2hs_{i}+\Delta E^{\rm{Q,eff}}_{i},$ (S31)
where
$\Delta E^{\rm{Q,eff}}_{i}=(\Delta\sigma)s_{i}\sum_{j\in{\mathcal{N}}_{i}}s_{j},$ (S32)
where $\Delta\sigma$ is evaluated at $T=0$, as shown in the inset of Fig. 2(a) of the main text. This effectively replaces the fermion diagonalization step with an effective correction from the fermions. In Fig. S7(b), we show the effective MC dynamics with the same initial random seed and same initial configurations as shown previously using exact MC. We next estimate the validity of this step by comparing $\Delta E_{i}^{\rm{Q}}$ from each Monte Carlo dynamics (see Eq. (S29)) with effective $\Delta E_{i}^{\rm{Q,eff}}$. The energy correction can be estimated for the two processes $s_{i}=1\rightarrow-1$ and for $s_{i}=-1\rightarrow 1$ for different configurations of neighborhood spins characterized by $N_{i}=\sum_{j\in{\mathcal{N}}_{i}}s_{j}$ that can take values $-4,-2,0,2,4$. In Fig. S7(c,d), we show the variation of $\Delta E^{\rm{Q,eff}}_{i}$ and $\Delta E_{i}^{\rm{Q}}$ for both the processes for different values of $N_{i}$ at $\kappa=1.2$. The values are essentially indistinguishable, establishing the validity and utility of the effective MC procedure. Thus, while exact droplet growth can have minor fluctuations between exact MC dynamics and effective MC dynamics, the effective MC method is particularly apt in evaluating the coarse-grained description of the system.
V.3 Typical and averaged droplet growth
Figure S8: Local magnetization: Local magnetization $s_{x,y}$ for unit cell at position $(x,y)$ where initial cluster-size with up-spins are $n=441$ ($n>n_{c}^{\rm{Q}}$), $n=225$ ($n_{c}^{\rm{Q}}>n>n_{c}^{\rm{cl}}$), $n=81$ ($n_{c}^{\rm{cl}}>n$) at times $t=0$, $t=20$, $t=100$ (in units of Monte Carlo steps per site) when $\kappa=0$ and $\kappa=1.2$. The parameters chosen are: $J=0.2$, $T=1/\beta=0.33$ (where $T_{c}\approx 0.45$), $h=0.02$. Number of unit cells in square lattice is $L^{2}=4096$.
Using the Monte Carlo simulation scheme described above, we compute local magnetization $s_{x,y}$ at location $(x,y)$ and time $t$ (in units of Monte Carlo steps per site or after $tL^{2}$ Monte Carlo steps where $L^{2}$ is the number of spins in square lattice) in a typical realization of the simulation. $s_{x,y}$ for various times $t$ in a typical realization for classical ($\kappa=0$) and topological ($20<|\kappa|<$ 2) situations are depicted in Fig. S8 where the boundary of the droplets can be clearly seen during their growth (shrinkage). As expected, any typical droplet boundary follows a very rugged trajectory.
To estimate the universal behavior of such growth, configuration averaging is performed which has been reported in the main text. An average local magnetization $\langle s_{x,y}\rangle$ where averaging is performed over $10^{3}$ realizations of the simulations is provided in Fig. 3(b) of the main text. Averaging over various realizations in $\langle s_{x,y}\rangle$ renders the growing (shrinking) droplets circularly symmetric.
VI Role of Slow Quench
Figure S9: Dynamics of average local magnetization under slow quench: (a) Linear quench protocol of $h(t)$ as a function of $t$ (see Eq. (S33)), (b,c) Average local magnetization $\langle s_{x,y}\rangle$ for the unit cell with coordinate $(x,y)$ for a cluster of up-spins having initial size $n=441$ at time $t=100$ (in the units of Monte Carlo steps per site) for different values of $\tau$ when (b) $\kappa=0$ and (c) $\kappa=1.2$. In all the plots, the parameters chosen are: $J=0.2$, $T=1/\beta=0.33$ (where $T_{c}\approx 0.45$), $h_{f}=0.02$. The number of unit cells in square lattice is $L^{2}=4096$ and number of simulations considered is $1000$. Cluster size at fixed $t=100$ reduces with increasing $\tau$.
In the main text we had shown the results for a sudden quench. We now analyze the growth and shrinkage of clusters under slow quench protocol of the external magnetic field $h(t)$ where we consider
$h(t)=\begin{cases}-h_{f}+\frac{2h_{f}t}{\tau},\text{~{}~{}~{}~{}for $0<t<\tau$,}\\ ~{}~{}~{}h_{f},\text{~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}for $t\geq\tau$.}\end{cases}$ (S33)
Thus, $\tau$ is the timescale of the quench (see Fig. S9(a)) over which $h(t)$ changes from $-h_{f}$ to $h_{f}$ (where we assume $h_{f}>0$) and the limit $\tau=0$ corresponds to the sudden quench considered in the main text. If $t$ and $\tau$ are chosen in the units of Monte Carlo steps per site, then we find that the size of a cluster having up-spins at a fixed time $t>\tau$ decreases with the increase of $\tau$ for both classical ($\kappa=0$) and topological ($20<|\kappa|<$ 2) situations, as can be seen from average local magnetization $\langle s_{x,y}\rangle$ in Fig. S9(b,c). This observation can be explained as follows: in this slow quench protocol, when $t<\tau/2$, $h(t)<0$ and thus a cluster having up-spins is not in the global free-energy minima and reduction of its size (i.e. growth of its larger background region) is more favored. Later, when $h(t)>0$ (for $t>\tau/2$), the cluster having up-spins reaches the global free-energy minima and starts showing growth (shrinkage) determined by usual nucleation physics. However, due to the reduction of its size when $t<\tau/2$, the size of a cluster at a fixed time $t>\tau$ becomes smaller compared to that for sudden quench situation ($\tau=0$). Thus, if a cluster grows under a sudden quench, it can be made to shrink when the quench protocol is changed to a slow quench performed over a time $\tau$, since the nucleation phenomenon is decided by the adiabatic limit where the instantaneous free-energy landscape determines the nucleation process. This leads to an increase in the effective critical cluster size with increasing $\tau$ in the slow quench protocol. Thus, the timescale $\tau$ can effectively tune the critical cluster size and influence the growth (shrinkage) of clusters.